• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name W. Deconinck

    Log entry time 20:07:40 on October 23, 2009

    Entry number 296359


    keyword=Overview of work performed in the hall earlier (with Ed, Alex)

    While in the hall we replace the He level units on the left arm Q2 and
    the right arm Q2 and Q3. When power cycling they gave repeatedly
    different values (off by 15%) and different from values on a known good
    unit (off by 60% in one case). The new units gave reasonable values of
    80% for all levels. The suspicion is that radiation damage (units
    located in the crawl space underneath the detector hut, open to lots of
    particle spraying) caused the units to become unstable and the power
    failure then gave them the last blow.

    After replacing the units the spectrometer seemed to come up fine, but
    unfortunately tripped soon afterward (see other entry, sorry that they
    are out of order, I am trying to keep up).